# Key varieties for prime $\mathbb{Q}$-Fano threefolds defined by Freudenthal triple systems

### v2

@@ 5.1. Definition
We note that the condition on $\mathfrak{S}_{\mathbb{A}}^{8}$ is equivalent to that $M_{x}$ and $M_{y}$ as in Subsection 4.1 have the same form.
Now we introduce different notation for coordinates of $\mathfrak{S}_{\mathbb{A}}^{8}$ as follows, with which we derive another presentation of the equations of $\mathfrak{S}_{\mathbb{A}}^{8}$ (Proposition 5.3) such that an group action on $\mathfrak{S}_{\mathbb{A}}^{8}$ will be more visible (Subsection 5.2):
@@
$V:=\left(\begin{array}[]{ccc}v_{2}&v_{1}&v_{0}\\ -v_{3}&-v_{2}&-v_{1}\end{array}\right):=\left(\begin{array}[]{ccc}y_{3}&x_{3}&t\\ -s&-y_{3}&-x_{3}\end{array}\right),$
@@ 5.3. $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$-fibration
Let $\mathsf{p}$ be a point of $\mathbb{A}_{D}^{4}$ and $F_{\mathsf{p}}$ the $\rho_{\mathfrak{S}}$-fiber over $\mathsf{p}.$ We use the descriptions of the $({\rm SL_{2}^{I}}\times(\mathbb{C}^{*})^{2})$-action on $\mathbb{A}_{D}^{4}$ as in Proposition 5.5 $(a)$$(d)$.
@@ 5.4. Summary of properties of $\mathfrak{S}_{\mathbb{A}}^{8}$
    The weighted projectivization $\mathfrak{S}_{\mathbb{P}}^{7}$ of $\mathfrak{S}_{\mathbb{A}}^{8}$ with some positive weights of coordinates is $\mathbb{Q}$-factorial and has Picard number one (we refer to Example 7.9 for an example of such a set of weights).
@@
where $A^{\dagger}$ and $Q^{\dagger}$ are the adjoint matrices of $A$ and $Q$, respectively, and $A,y_{1},y_{2}$, $\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ x_{3}\end{array}\right),\left(\begin{array}[]{c}x_{4}\\ x_{5}\\ x_{6}\end{array}\right)$ in ibid. correspond to $Q,-t,s,x,y$ here, respectively. For, we can verify that
@@ 2. $(2)$
(1). We can check that the $x_{i}$-, $y_{i}$-, $s$-, and $t$-charts ($i=1,2,3)$ of $\mathfrak{Z}_{\mathbb{A}}^{12}$ are smooth and the union of them is irreducible and is of codimension $4$ in the ambient affine space. Using [DGPS], we can show that the ideal of $\mathfrak{Z}_{\mathbb{A}}^{12}$ has the $9\times 16$ graded minimal free resolution of length $4$. Therefore, by the proofs of [Tak6, Prop. 4.8], we obtain (1).
(2). We can check that $(R_{\mathfrak{Z}})_{x_{1}}$ is a localization of a polynomial ring. We can also check that $\mathfrak{Z}_{\mathbb{A}}^{12}\cap\{x_{1}=0\}$ is normal. Indeed, on the $x_{2}$-, $x_{3}$-, $y_{2}$- and $y_{3}$-charts, $\mathfrak{Z}_{\mathbb{A}}^{12}\cap\{x_{1}=0\}$ is smooth, and $\mathfrak{Z}_{\mathbb{A}}^{12}\cap\{x_{1}=x_{2}=x_{3}=y_{2}=y_{3}=0\}$ is of codimension $2$ in $\mathfrak{Z}_{\mathbb{A}}^{12}\cap\{x_{1}=0\}$. Therefore The proofs of [Tak6, Prop. 4.9] work verbatim for (2). ∎
The main purpose of this section is to interpret $\mathfrak{Z}_{\mathbb{A}}^{12}$ in the context of [Tak5] as follows:
@@
We define $\mathfrak{Z}_{\mathbb{P}}^{11}$ to be the projective variety obtained from $\mathfrak{Z}_{\mathbb{A}}^{12}$ by setting $w(s)=w(t)=2$ and all the other weights of coordinates as $1$ ( Note that $\mathfrak{Z}_{\mathbb{P}}^{11}$ is contained in $\mathbb{P}(1^{14},2^{2})$). The variety $\mathfrak{Z}_{\mathbb{P}}^{11}$ is isomorphic to the $\mathbb{Q}$-Fano variety $\Sigma$ as in [Tak5, Thm.1.1] associated to prime $\mathbb{Q}$-Fano $3$-folds of No.1.1 in [Tak1].
@@
Note that the projective variety ¯:=$\overline{\mathfrak{Z}}:=\left\{{}^{t}\!xPy=0,{}^{t}\!xQy=0\right\}\subset\mathbb{P}^{13}$ with $P=E$ and ${\rm Tr\,}Q=0$ can be identified with $\overline{\Sigma}$ as in [Tak5, Subsec.4.1.1] with the obvious correspondence between the coordinates of them.
By the construction of $\Sigma$ from $\overline{\Sigma}$ as in [Tak5, Sec.5], $\Sigma$ is isomorphic to $\overline{\Sigma}$ in codimension $1$ except the image $\overline{\Gamma}$ in $\overline{\Sigma}$ of the exceptional divisor for the blow-up of two $1/2(1^{11})$-singularities.
[…]
By Proposition 6.3 and [Tak5, Thm.1.1], we obtain the following result of Gushel´-Mukai type:
@@
Any prime $\mathbb{Q}$-Fano $3$-fold of No.1.1 in [Tak1] is a weighted complete intersection of $\mathfrak{Z}_{\mathbb{P}}^{11}$ as in Proposition 6.3 with respect to hypersurfaces of weight $1$.
Finally, we describe an ${\rm SL}_{3}$ on $\mathfrak{Z}_{\mathbb{A}}^{12}$ as follows:
@@
Let $\mathbb{A}_{\mathfrak{Z}}^{16}$ be the affine subspace $\{{\rm Tr}\,Q=0\}$ in the affine space whose coordinates are $s,\,t$ and the entries of $x,y,Q$. The affine space $\mathbb{A}_{\mathfrak{Z}}^{16}$ has the ${\rm SL}_{3}$-action preserving $\mathfrak{Z}_{\mathbb{A}}^{12}$ defined by
@@ 7.1. The $G_{2}^{(4)}$-cluster variety $\mathfrak{Cl}_{\mathbb{A}}^{10}$
In this paper, however, we call a smaller variety as the $G_{2}^{(4)}$-cluster variety. By the big table [Table], we observe that $A_{12}$ and $A_{34}$ is always nonzero constant when a $\mathbb{Q}$-Fano 3-fold is constructed from a weighted projectivization of the $G_{2}^{(4)}$-cluster variety. Then, replacing $A_{1}$, $A_{2}$, $A_{3}$, $A_{4}$, $\lambda_{13}$, $\lambda_{24}$ with $A_{12}^{-1}A_{1}$, $A_{12}^{-1}A_{2}$, $A_{34}^{-1}A_{3}$, $A_{34}^{-1}A_{4}$, $A_{12}^{-1}A_{34}^{-1}\lambda_{13}$, $A_{12}^{-1}A_{34}^{-1}\lambda_{24}$, we see that it is possible to set $A_{12}=A_{34}=1$.
@@
$x={}^{t}\!\left(\begin{array}[]{ccc}\theta_{4}&\theta_{1}&A_{23}\end{array}\right),y={}^{t}\!\left(\begin{array}[]{ccc}A_{41}&\theta_{2}&\theta_{3}\end{array}\right),$
@@
We can immediately check that this definition of the $G_{2}^{(4)}$-cluster variety coincides with that in [CD, Subset. 1.2.2] when $A_{12}=A_{34}=1$.
@@
All the equations of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ are weighted homogeneous if and only if the following conditions on the weights of coordinates hold:
@@
In the following subsections, we only consider the weights of coordinates as in Proposition 7.2 such that the weighted projectivizations of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ itself or its subvarieties associated to these produce prime $\mathbb{Q}$-Fano 3-folds. The list of such weights are presented in [Table].
@@ 7.2. The maximal case
By the big table [Table], we observe that all the weights of the coordinates of the $G_{2}^{(4)}$-cluster variety $\mathfrak{Cl}_{\mathbb{A}}^{10}$ is positive only in the two cases No.5530 and No.11455 of [GRDB]. We may verify the following by a straightforward calculation.
@@
The equations of the affine variety $\mathfrak{Cl}_{\mathbb{A}}^{10}$ is presented in the format of the equations of $\mathfrak{U}_{\mathbb{A}}^{14}$ by setting
$x={}^{t}\!\left(\begin{array}[]{ccc}\theta_{1}-A_{2}A_{23}&\theta_{4}-A_{3}A_{23}&A_{23}\end{array}\right),y={}^{t}\!\left(\begin{array}[]{ccc}A_{41}&1/2(\theta_{3}-\theta_{2})&1/2(\theta_{3}+\theta_{2})\end{array}\right),$
@@
By Proposition 7.2 and the weights of coordinates of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ for No.5530 and No.11455, we see that all the entries of $x,y,P,Q$ as in Proposition 7.3 are weighted homogeneous. Therefore we have the following:
@@
Moreover, the isomorphism is weighted homogeneous for No.5530 and No.11455 with the corresponding weights of coordinates of (7.1).
7.3. The case only one coordinate is a nonzero constant
By [Table], there are 12 classes of prime $\mathbb{Q}$-Fano 3-folds for which only one coordinate of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ (actually $A_{1}$,$A_{3}$ or $A_{4}$) is a nonzero constant. By symmetry, we may identify the three cases, so we have only to consider the case that $A_{4}$ is a nonzero constant; this is the case for No.1169, 1182, 5860, 5870, 16228. We denote by $\mathfrak{Cl}_{\mathbb{A},A_{4}}^{9}$ the subvariety of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ with $A_{4}=-1$. By suitable changes of scales of coordinates, we may assume that $A_{4}=-1$.
@@ 7.3. The case only one coordinate is a nonzero constant
The equations of the affine variety $\mathfrak{Cl}_{\mathbb{A},A_{4}}^{9}$ is presented in the format of the equations of $\mathfrak{U}_{\mathbb{A}}^{14}$ by setting
$x={}^{t}\!\left(\begin{array}[]{ccc}A_{41}&\theta_{2}&\theta_{3}\end{array}\right),y={}^{t}\!\left(\begin{array}[]{ccc}A_{23}&\theta_{1}&\theta_{4}-1/2\lambda_{24}A_{23}\end{array}\right),$
@@
Moreover, the isomorphism is weighted homogeneous with the corresponding weights of coordinates of (7.2).
In the following subsections, we consider the cases in which exactly two of the coordinates are nonzero constants. There are 4 cases; the set of the two nonzero constant coordinates is $\{A_{3},A_{4}\}$ , $\{A_{1},A_{4}\}$, $\{A_{2},A_{3}\}$, or $\{A_{1},A_{3}\}$. For them, we may assume that the nonzero constants are $-1$. We denote by $\mathfrak{Cl}_{\mathbb{A},A_{i}\!A_{j}}^{8}$ the subvariety of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ with $A_{i}=A_{j}=-1$. We will treat the 4 cases separately in the sequel.
7.4. The case $A_{3}$ and $A_{4}$ are nonzero constants

### v3

@@ 5.1. Definition
We recall that the matrices $M_{x}$ and $M_{y}$ are defined in Subsection 4.1. We note that the condition on $\mathfrak{S}_{\mathbb{A}}^{8}$ is equivalent to that $M_{x}$ coincides with $M_{y}$ by replacing $x_{1},x_{2},x_{3}$ with $y_{1},y_{2},y_{3}$ respectively.
Now we introduce a different notation for coordinates of $\mathfrak{S}_{\mathbb{A}}^{8}$ as follows, with which we derive another presentation of the equations of $\mathfrak{S}_{\mathbb{A}}^{8}$ (Proposition 5.3) such that a group action on $\mathfrak{S}_{\mathbb{A}}^{8}$ will be more visible (Subsection 5.2):
@@
$V:=\left(\begin{array}[]{ccc}v_{2}&v_{1}&v_{0}\\ -v_{3}&-v_{2}&-v_{1}\end{array}\right):=\left(\begin{array}[]{ccc}y_{3}&x_{3}&t\\ -s&-y_{3}&-x_{3}\end{array}\right).$
@@ 5.3. $\mathbb{P}^{1}\times\mathbb{P}^{1}\times\mathbb{P}^{1}$-fibration
Let $\mathsf{p}$ be a point of $\mathbb{A}_{D}^{4}$ and $F_{\mathsf{p}}$ the $\rho_{\mathfrak{S}}$-fiber over $\mathsf{p}.$ We use the descriptions of the $({\rm SL_{2}^{I}}\times(\mathbb{C}^{*})^{2})$-action on $\mathbb{A}_{D}^{4}$ as in Proposition 5.5 (a)(a)$(d)$.
@@ 5.4. Summary of properties of $\mathfrak{S}_{\mathbb{A}}^{8}$
    The weighted projectivization $\mathfrak{S}_{\mathbb{P}}^{7}$ of $\mathfrak{S}_{\mathbb{A}}^{8}$ with some positive weights on the coordinates is $\mathbb{Q}$-factorial and has Picard number one (we refer to Example 7.9 for an example of such a set of weights).
@@
where $A^{\dagger}$ and $Q^{\dagger}$ are the adjoint matrices of $A$ and $Q$, respectively, and $A,y_{1},y_{2}$, $\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ x_{3}\end{array}\right),\left(\begin{array}[]{c}x_{4}\\ x_{5}\\ x_{6}\end{array}\right)$ in ibid. correspond to Q,t,s,x,yQ,-t,s,x,y here, respectively. We can verify that
@@ 2. $(2)$
(1). We can check that the $x_{i}$-, $y_{i}$-, $s$-, and $t$-charts ($i=1,2,3)$ of $\mathfrak{Z}_{\mathbb{A}}^{12}$ are smooth and the union of them is irreducible and is of codimension $4$ in the ambient affine space. Using [DGPS], we can show that the ideal of $\mathfrak{Z}_{\mathbb{A}}^{12}$ has the $9\times 16$ graded minimal free resolution of length $4$. Therefore, by the proofs of [Tak4, Prop.4.8], we obtain (1).
(2). We can check that $(R_{\mathfrak{Z}})_{x_{1}}$ is a localization of a polynomial ring. We can also check that $\mathfrak{Z}_{\mathbb{A}}^{12}\cap\{x_{1}=0\}$ is normal. Indeed, on the $x_{2}$-, $x_{3}$-, $y_{2}$- and $y_{3}$-charts, $\mathfrak{Z}_{\mathbb{A}}^{12}\cap\{x_{1}=0\}$ is smooth, and $\mathfrak{Z}_{\mathbb{A}}^{12}\cap\{x_{1}=x_{2}=x_{3}=y_{2}=y_{3}=0\}$ is of codimension $2$ in $\mathfrak{Z}_{\mathbb{A}}^{12}\cap\{x_{1}=0\}$. Therefore The proofs of [Tak4, Prop.4.9] work verbatim for (2). ∎
The main purpose of this section is to interpret $\mathfrak{Z}_{\mathbb{A}}^{12}$ in the context of [Tak3] as follows:
@@
We define $\mathfrak{Z}_{\mathbb{P}}^{11}$ to be the projective variety obtained from $\mathfrak{Z}_{\mathbb{A}}^{12}$ by setting $w(s)=w(t)=2$ and all the other weights on the coordinates as 11 ( Note that $\mathfrak{Z}_{\mathbb{P}}^{11}$ is contained in $\mathbb{P}(1^{14},2^{2})$). The variety $\mathfrak{Z}_{\mathbb{P}}^{11}$ is isomorphic to the $\mathbb{Q}$-Fano variety $\Sigma$ as in [Tak3, Thm.1.1] associated to prime $\mathbb{Q}$-Fano $3$-folds of No.1.1 in [Tak1].
@@
Note that the projective variety $\overline{\mathfrak{Z}}:=\left\{{}^{t}\!xPy=0,{}^{t}\!xQy=0\right\}\subset\mathbb{P}^{13}$ with $P=E$ and ${\rm Tr\,}Q=0$ can be identified with $\overline{\Sigma}$ as in [Tak3, Subsec.4.1.1] with the obvious correspondence between their coordinates.
By the construction of $\Sigma$ from $\overline{\Sigma}$ as in [Tak3, Sec.5], $\Sigma$ is isomorphic to $\overline{\Sigma}$ in codimension $1$ except the image $\overline{\Gamma}$ in $\overline{\Sigma}$ of the exceptional divisor for the blow-up of two $1/2(1^{11})$-singularities.
We show that a similar fact holds for $\mathfrak{Z}_{\mathbb{P}}^{11}$ and $\overline{\mathfrak{Z}}$. We consider the rational map $\pi\colon\mathfrak{Z}_{\mathbb{P}}^{11}\dashrightarrow\mathfrak$$\overline${Z}} which is the restriction of the projection from the $(s:t)$-line. Note that $\pi$ is defined on $U:=\mathfrak{Z}_{\mathbb{P}}^{11}\setminus\{\text{the\,}s\text{-, $t$-points\}. }$ We set $\overline{\Delta}:=\left\{x=0,y^{\sharp}=0\right\}\cup\left\{y=0,x^{\sharp}=0\right\}\subset\overline{\mathfrak{Z}}.$ From the equation of $\mathfrak{Z}_{\mathbb{P}}^{11}$, we see that the closure of the inverse image of $\overline{\Delta}|_{\pi(U)}$ by $\pi|_{U}$ is $\Delta:=\left\{x=0,y^{\sharp}=0,t=0\right\}\cup\left\{y=0,x^{\sharp}=0,s=0\right\}.$ By the equations of $\mathfrak{Z}_{\mathbb{P}}^{11}$ and $\overline{\mathfrak{Z}}$, we see that, if $x\not=0$ and $y\not=0$ on $\overline{\mathfrak{Z}}$, then both $s$ and $t$ are recovered by the equations of $\mathfrak{Z}_{\mathbb{P}}^{11}$. Moreover, we see that points of $\{x=0\}\cup\{y=0$} outside $\overline{\Delta}$ are not the $\pi$-images of points of $U$, and the fiber of $\pi|_{U}$ over a point ${\rm p}\in\overline{\Delta}$ is $1$-dimensional if ${\rm p}\not\in\{x=y=0\}$, and $2$-dimensional if ${\rm p}\in\{x=y=0\}$. Therefore the $\pi|_{U}$-exceptional locus is $\Delta|_{U}$. By the equation of $\mathfrak{Z}_{\mathbb{P}}^{11}$ and [Tak3, Prop.4.4], the locus $\overline{\Delta}$ coincides with $\rm Sing$$\mathfrak$$\overline${Z}}, which is of codimension 3 in $\overline{\mathfrak{Z}}$. Therefore the exceptional locus of $\pi|_{U}$ is of codimension 2 in $\mathfrak{Z}_{\mathbb{P}}^{11}$, which implies $\mathfrak{Z}_{\mathbb{P}}^{11}$ is isomorphic to $\overline{\mathfrak{Z}}$ in codimension $1$ except $\{x=0\}\cup\{y=0$}. Note that $\{x=0\}\cup\{y=0$} is identified with $\overline{\Gamma}$. Therefore, we see that $\Sigma$ and $\mathfrak{Z}_{\mathbb{P}}^{11}$ are isomorphic in codimension $1$ identifying $\overline{\Sigma}$ and $\overline{\mathfrak{Z}}$. By [Tak3, Thm.1.1], $\Sigma$ is a $\mathbb{Q}$-factorial $\mathbb{Q}$-Fano variety with Picard numbers 1, and, by Lemma 6.2 and the proof of [Tak4, Cor.4.10], so is $\mathfrak{Z}_{\mathbb{P}}^{11}$. Therefore, $\Sigma$ and $\mathfrak{Z}_{\mathbb{P}}^{11}$ are actually isomorphic by [Tak2, Lem.5.5] as desired. ∎
By Proposition 6.3 and [Tak3, Thm.1.1], we obtain the following result of Gushel-Mukai type:
@@
Any prime $\mathbb{Q}$-Fano $3$-fold of No.1.1 in [Tak1] is a weighted complete intersection of $\mathfrak{Z}_{\mathbb{P}}^{11}$ as in Proposition 6.3 with respect to hypersurfaces of weight 11.
Finally, we describe an ${\rm SL}_{3}$-action on $\mathfrak{Z}_{\mathbb{A}}^{12}$ as follows:
@@
Let $\mathbb{A}_{\mathfrak{Z}}^{16}$ be the affine subspace $\{{\rm Tr}\,Q=0\}$ in the affine space whose coordinates are s,ts,\,t and the entries of $x,y,Q$. The affine space $\mathbb{A}_{\mathfrak{Z}}^{16}$ has the ${\rm SL}_{3}$-action preserving $\mathfrak{Z}_{\mathbb{A}}^{12}$ defined by
@@ 7.1. The $G_{2}^{(4)}$-cluster variety $\mathfrak{Cl}_{\mathbb{A}}^{10}$
In this paper, however, we call a smaller variety as the $G_{2}^{(4)}$-cluster variety. By the big table [Table], we observe that $A_{12}$ and $A_{34}$ are always nonzero constants when a $\mathbb{Q}$-Fano 3-fold is constructed from a weighted projectivization of the $G_{2}^{(4)}$-cluster variety. Then, replacing $A_{1}$, $A_{2}$, $A_{3}$, $A_{4}$, $\lambda_{13}$, $\lambda_{24}$ with $A_{12}^{-1}A_{1}$, $A_{12}^{-1}A_{2}$, $A_{34}^{-1}A_{3}$, $A_{34}^{-1}A_{4}$, $A_{12}^{-1}A_{34}^{-1}\lambda_{13}$, $A_{12}^{-1}A_{34}^{-1}\lambda_{24}$, we see that it is possible to set $A_{12}=A_{34}=1$.
@@
[…]
@@
We can immediately check that this definition of the $G_{2}^{(4)}$-cluster variety coincides with that in [CD, Subsec.1.2.2] when $A_{12}=A_{34}=1$.
@@
All the equations of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ are weighted homogeneous if and only if the following conditions for the weights on the coordinates hold:
@@
In the following subsections, we only consider the weights on the coordinates as in Proposition 7.2 such that the weighted projectivizations of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ itself or its subvarieties associated to these produce prime $\mathbb{Q}$-Fano 3-folds. The list of such weights are presented in [Table].
@@ 7.2. The maximal case
By the big table [Table], we observe that all the weights on the coordinates of the $G_{2}^{(4)}$-cluster variety $\mathfrak{Cl}_{\mathbb{A}}^{10}$ are positive only in the two cases No.5530 and No.11455 of [GRDB]. We may verify the following by a straightforward calculation.
@@
The equations of the affine variety $\mathfrak{Cl}_{\mathbb{A}}^{10}$ are presented in the format of the equations of $\mathfrak{U}_{\mathbb{A}}^{14}$ by setting
[…]
@@
By Proposition 7.2 and the weights on the coordinates of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ for No.5530 and No.11455, we see that all the entries of $x,y,P,Q$ as in Proposition 7.3 are weighted homogeneous. Therefore we have the following:
@@
Moreover, the isomorphism is weighted homogeneous for No.5530 and No.11455 with the corresponding weights on the coordinates of (7.1).
7.3. The case where only one coordinate is a nonzero constant
By [Table], there are 12 classes of prime $\mathbb{Q}$-Fano 3-folds for which only one coordinate of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ (actually $A_{1}$$A_{3}$ or $A_{4}$) is a nonzero constant. By symmetry, we may identify the three cases, so we have only to consider the case that $A_{4}$ is a nonzero constant; this is the case for No.1169, 1182, 5860, 5870, 16228. We denote by $\mathfrak{Cl}_{\mathbb{A},A_{4}}^{9}$ the subvariety of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ with $A_{4}=-1$. By suitable changes of scales of coordinates, we may assume that $A_{4}=-1$.
@@ 7.3. The case where only one coordinate is a nonzero constant
The equations of the affine variety $\mathfrak{Cl}_{\mathbb{A},A_{4}}^{9}$ are presented in the format of the equations of $\mathfrak{U}_{\mathbb{A}}^{14}$ by setting
[…]
@@
Moreover, the isomorphism is weighted homogeneous with the corresponding weights on the coordinates of (7.2).
In the following subsections, we consider the cases in which exactly two of the coordinates are nonzero constants. There are 4 cases; the set of the two nonzero constant coordinates is $\{A_{3},A_{4}\}$ , $\{A_{1},A_{4}\}$, $\{A_{2},A_{3}\}$, or $\{A_{1},A_{3}\}$. For them, we may assume that the nonzero constants are 1-1. We denote by $\mathfrak{Cl}_{\mathbb{A},A_{i}\!A_{j}}^{8}$ the subvariety of $\mathfrak{Cl}_{\mathbb{A}}^{10}$ with $A_{i}=A_{j}=-1$. We will treat the 4 cases separately in the sequel.
7.4. The case where $A_{3}$ and $A_{4}$ are nonzero constants